OpenStudy (anonymous):

Find the number of identical terms in the following two series: a) 3,7,11,15,...,367 b) 2,9,16,23,...,709

OpenStudy (amistre64):

find the intersection of 2 lines ...

OpenStudy (amistre64):

y = 3+4n y = 2+7n

OpenStudy (anonymous):

what's the answer?

OpenStudy (amistre64):

thats not a very good question ...

OpenStudy (amistre64):

we are not a "whats the answer" website

OpenStudy (anonymous):

i mean that kindly give me a solution

OpenStudy (amistre64):

i have given you information which you should be capable of working out .. if you have questions about what ive offered, then as for clarification .. not for "the answer". We can work towards the solution, but I need you to give me something to work with.

OpenStudy (amistre64):

the common terms will represent the intersection of the respective rules of each .. and they are not series, they are sequences. a series is a sum of a sequence, whereas a sequence is a list of terms seperated by commas

OpenStudy (anonymous):

thanks...is the answer 102 if i did it right?

OpenStudy (anonymous):

no.. is the answer 13

OpenStudy (amistre64):

hmm, i have to admit that they will not have the same nth term , so my original idea, no matter how spectaular ...is flawed lol

OpenStudy (anonymous):

hey... i found that it's an A.P. and then i got answer 13

OpenStudy (amistre64):

AP is correct, 13 may be one value they have in common, but is it the only value they share?

OpenStudy (anonymous):

other value i got is 102...but it's not possible...as sequence b has 102 terms

OpenStudy (amistre64):

i cant get a 13, canyou show me how you got to that?

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

a) 3,7,11,15,...,367 ^^ 13 is missing b) 2,9,16,23,...,709 ^^ 13 is missing

OpenStudy (amistre64):

by trial and error, i get 51 to be a common term between them

OpenStudy (anonymous):

sequence 'a' has 92 terms and 'b' has 102 let pth term of a =qth term of b 3+(p-1)4=2+(q-1)7 4p-1=7q-5 4(p+1)=7q (p+1)/7=q/4=k (say) so p=7k-1 and q=4k now p<=92 and q<=102 so k<=93/7 and k<=51/2 so k<=13+2/7 so k=13

OpenStudy (anonymous):

can there be a quicker way to solve this?

OpenStudy (amistre64):

that looks similar to what i just developed on the paper in front if me :) let a = 3+4k , for some integer k let b = 2+7n , for some integer n when a=b we have: 3+4k = 2+7n or written another way 4n - 7k = -1

OpenStudy (amistre64):

we can either view this as a euler number theory problem, or simply as a line that we can plot and compare integer solutions with

OpenStudy (amistre64):

if we do a euler run on it; we need to find the gcd of 4 and 7 ... 7 = 4(1) + 3 4 = 3(1) + 1 <-- well, the gcd is already 1 so we dont have to modify it any

OpenStudy (amistre64):

maybe modify it by a negative :) does this look like the content of your material? or shold we approach it another way?

OpenStudy (anonymous):

phew... finally got my internet working

OpenStudy (amistre64):

if we use a line approach: 4n - 7k = -1 , we can define n in terms of k n = (7k-1)/4 k=0, n=-1/4 not an integer k=1, n=6/4 not an integer k=2, n=13/4 not an integer k=3, n=20/4 = 5, thats good k=4,5,6, n is not an integer k=7, n=48/4 = 12 , thats good seems to me that every 4th k produces an integer result so when k = 3+4t, for t=0,1,2,3,4,.... n = (7(3+4t)-1)/4 n = (21+28t -1)/4 n = (20+28t)/4 n = 5+7t

OpenStudy (amistre64):

the question is now, how many ts can we fit into 91 and 101 terms?

OpenStudy (amistre64):

might be a simpler approach 3+4k = 367 ; k = 91 91 /4 = 22.75 so im thinking there are at best 22 or 23 terms they have in common

OpenStudy (amistre64):

im really hoping some if this makes sense :)

OpenStudy (anonymous):

any way to check it?

OpenStudy (amistre64):

make a tables in excel and compare it side by side

OpenStudy (anonymous):

maybe i'll try a java program

OpenStudy (amistre64):

there are 25 as is, but i havent accounted for the limit of n and k in the number of terms

OpenStudy (amistre64):

k can at best go to 101 ... n=55, k=96 is the best we can accomplish with this, and that adds up to 14 terms for me

OpenStudy (anonymous):

maybe taking it simply as an AP gives 13, and maybe that's fine

OpenStudy (amistre64):

i think i have my n and k bacwards ... let me rechk

OpenStudy (amistre64):

yeah, i had them backwards :/ k on the left, n on the right, i got to 13 eventually:)

OpenStudy (anonymous):

so 13 is alright THANKS amistre64

OpenStudy (amistre64):

yep, let me see if i can make that word out a little better :) condense the beast! let a = 3+4k , for some k = 0,1,2,3, ... , 3+4k = 367 ; 91 let b = 2+7n , for some n = 0,1,2,3,..., 2+7n = 709 ; 101 we can work it up to a setup for some t = 0,1,2,3,4,... k = 3+4t 91 = 3+4t 22 = t n = 5+7t 101 = 5+7t 13.714 = t notice that this is only valid for the smaller value of t, so when t = 13

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

good luck :)